[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Aug 18, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Aug 19 02:35:51 UTC 2017


HSU Shift Log for Fri Aug 18 2017 16:40:29 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Brian replaced graphics card on Tuna. Encountered some  
problems getting the machine to boot back up, but seems to be stable now.

AWIPS-2/NCP issues: None

Other Techincal issues: Atlantic/Pacific desk phones aren't  
forwarding/being forwarded to properly, so all phones don't always ring.  
NCO ticket opened. Brief GOES-16 outage from 1145-1230 UTC. Resolved by  
Monica/Chris M.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170819/1eb120d7/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list